Find at least one solution to the following equation:
sin(x^2 − 1)/1 − sin(x^2 − 1) = sin(x) + sin^2(x) + sin^3(x) + sin^4(x) + · · ·
1
Expert's answer
2020-11-17T16:39:00-0500

Firstly, let us solve the equation "x^2-1=x" which is equivalent to "x^2-x-1=0", and has the solutions "a=\frac{1+\sqrt{5}}{2}" and "b=\frac{1-\sqrt{5}}{2}". Therefore, "\sin(a^2-1)=\sin(a)" and "\sin(b^2-1)=\sin(b)". Since "\frac{\pi}{2}<a<\pi" and "-\frac{\pi}{2}<b<0", we conclude that "|\sin a|<1" and "|\sin b|<1". Then the

infinite geometric progression "\sin(a), \sin^2(a), \sin^3(a),..., \sin^4(a), ..." has a scale factor "\sin a" and the common ratio "\sin a" with "|\sin a|<1". Consequently,

"\sin(a) + \sin^2(a) + \sin^3(a) + \sin^4(a) + · · ·=\frac{\sin(a)}{1-\sin(a)}".

By analogy, "\sin(b) + \sin^2(b) + \sin^3(b) + \sin^4(b) + · · ·=\frac{\sin(b)}{1-\sin(b)}".

Therefore, "\frac{\sin(a^2-1)}{1-\sin(a^2-1)}=\frac{\sin(a)}{1-\sin(a)}" and "\frac{\sin(b^2-1)}{1-\sin(b^2-1)}=\frac{\sin(b)}{1-\sin(b)}", and we conclude that "a" and "b" are the solutions of the equation "\frac{\sin(x^2 − 1)}{1 − \sin(x^2 − 1)} = \sin(x) + \sin^2(x) + \sin^3(x) + \sin^4(x) + · · ·"